# IPO: Iterative Preference Optimization for Text-to-Video Generation

### v2

@@ Abstract
Video foundation models have achieved significant advancement with the help of network upgrade as well as model scale-up. However, they are still hard to meet requirements of applications due to unsatisfied generation quality. To solve this problem, we propose to align video foundation models with human preferences from the perspective of post-training in this paper. Consequently, we introduce an Iterative Preference Optimization strategy to enhance generated video quality by incorporating human feedback. Specifically, IPO exploits a critic model to justify video generations for pairwise ranking as in Direct Preference Optimization or point-wise scoring as in Kahneman-Tversky Optimization. Given this, IPO optimizes video foundation models with guidance of signals from preference feedback, which helps improve generated video quality in subject consistency, motion smoothness and aesthetic quality, etc. In addition, IPO incorporates the critic model with the multi-modality large language model, which enables it to automatically assign preference labels without need of retraining or relabeling. In this way, IPO can efficiently perform multi-round preference optimization in an iterative manner, without the need of tediously manual labeling. Comprehensive experiments demonstrate that the proposed IPO can effectively improve the video generation quality of a pretrained model and help a model with only 2B parameters surpass the one with 5B parameters. Besides, IPO achieves new state-of-the-art performance on VBench benchmark. We will release our source codes, models as well as dataset to advance future research and applications.
Figure 1: The overview of iterative preference optimization(IPO) framework
@@ 1 Introduction
Thus, we propose to align Text-to-Video models (T2Vs) with human preference from the perspective of post-training in this paper. In particular, we present the Iterative Preference Optimization (IPO) framework for enhancing T2Vs. IPO introduces preference priors for finetuning T2Vs in the form of reinforcement learning, which considers subject consistency, motion smoothness and aesthetic quality, etc. This leads to improved video generations that are well aligned with human preference. In addition, difference from prior single-round methods, IPO adopts a multi-round optimization strategy to reinforce T2Vs in an iterative manner, which strengths the models continuously. In this way, IPO can effectively improve T2Vs without need of large-scale dataset and expensive supervised finetuning.
@@ 2 Related Work
Recent advances in video generation have been predominantly driven by diffusion models [16, 40, 41, 55], which have demonstrated remarkable improvements across multiple dimensions, including diversity, fidelity, and image quality [5, 7, 17, 26, 38, 47, 49], with the scaling of training data and model size further enhancing their performance [14, 39, 48]. In the domain of Text-to-Video (T2V) generation, pre-training models from scratch demands substantial computational resources and access to extensive human-curated video datasets, highlighting the critical importance of leveraging large-scale datasets for effective model training [10, 29, 20, 3, 19]. Current methodologies, such as those presented in [2, 46, 12], build upon text-to-image models by incorporating spatial attention mechanisms, introducing temporal attention modules, and fine-tuning them on video datasets, often employing joint image and video training strategies [30], with [12] proposing a plug-and-play temporal module that facilitates video generation from personalized image models. State-of-the-art results in terms of pixel quality and temporal consistency have been achieved by methods such as [56, 4], while recent diffusion models based on the Dit structure, such as [3, 52, 4], have demonstrated exceptional performance, advancing the frontiers of T2V generation by enhancing spatiotemporal coherence and scalability, showcasing the potential of Transformer-based diffusion architectures. Further advancements have been made in addressing the discrepancy between spatial and temporal modules in two-stage training, with works such as [24, 19, 52] employing 3D full attention mechanisms to improve overall performance and effectively bridge the gap between spatial and temporal components.
The training of large models typically involves two key stages: pre-training and post-training, with relatively limited research focusing on the post-training phase for text-to-video (T2V) generation. In this context, [53] introduced a temporal decaying reward (TAR) mechanism, grounded in the assumption that the central frame of a video should be assigned the highest importance, while the emphasis gradually diminishes towards the peripheral frames. This strategic allocation of importance across frames ensures a more stable and visually coherent video generation process. Meanwhile, [21] explored the extensive design space of energy functions to enhance ordinary differential equation (ODE) solvers, demonstrating its potential by improving T2V model training through the extraction of motion priors from training videos. Additionally, VADER [35] proposed a method that leverages the gradient of a reward model to fine-tune a base video diffusion model, utilizing a wide range of pre-trained vision models to align various video diffusion models. This alignment approach, facilitated by VADER, exhibits strong generalization capabilities, even for cues not encountered during training, thereby advancing the adaptability and performance of T2V systems.
Aligning with human preferences
Aligning human preferences through post-training has been widely studied for large language models (LLMs), with several approaches [31, 37, 59] making notable advancements. Reinforcement Learning from Human Feedback (RLHF) [31] trains a reward function from comparison data on model outputs to align the policy model. Rank Responses for Human Feedback (RRHF) [54], a simpler alternative to Proximal Policy Optimization (PPO), extends supervised fine-tuning and reward model training. Sequence Likelihood Calibration (SLiC) [58] achieves alignment using human feedback data collected for different models, akin to offline RL data. Direct Preference Optimization (DPO) [36] simplifies alignment by directly training on human preference data, improving stability and performance while reducing computational costs. Kahneman-Tversky Optimization (KTO) [8] maximizes utility based on the Kahneman-Tversky model, rather than maximizing log-likelihood of preferences. Rejection Sampling Optimization (RSO) [28] introduces statistical rejection sampling to derive preference data from the optimal policy, enhancing alignment with human preferences.
Several methods have successfully integrated preference alignment with diffusion models, driving significant progress in the field. DraFT [6] fine-tunes models to maximize differentiable reward functions, such as human preference scores, while [34] aligns models with downstream reward functions via end-to-end backpropagation through the denoising process. Similarly, [50] optimizes models based on reward scores, but these approaches often face inefficiency and instability. To overcome these challenges, DPOK [9] frames fine-tuning as a reinforcement learning problem, combining policy optimization with KL regularization and policy gradients to update pre-trained text-to-image diffusion models. DDPO [1] reframes denoising as a multi-step decision problem, using policy gradients to enhance cue-image alignment without additional data collection or human annotation. Diffusion-DPO [45] optimizes models using human comparison data, reformulating Direct Preference Optimization (DPO) to improve model likelihood and differentiable objectives. Diffusion-KTO [22] introduces a novel approach for aligning text-to-image models by maximizing expected human utility, removing the need for costly pairwise preference data. The iterative preference optimization algorithm in this work further enhances offline learning methods like Diffusion-DPO and Diffusion-KTO, addressing instability and performance issues while advancing alignment effectiveness.
@@ 3 METHOD
In this section, we present our proposed general iterative preference optimization (IPO) framework, which is designed to support both pairwise and pointwise preference alignment optimization algorithms. For offline preference alignment algorithms, the effectiveness is often hindered by the limitations of offline datasets, as highlighted by RSO [27], particularly when the offline data distribution diverges from the target distribution. Taking Direct Preference Optimization (DPO) [36] as a representative example, which can be extended to other algorithms, DPO can be interpreted as imposing constraints on the resulting policy through the dataset $D$. Informally, for DPO to converge to the optimal policy $\pi^{*}$, it requires an infinite dataset $D$ to comprehensively cover the entire query-response space. However, in practice, training on a finite dataset $D$ inevitably fails to encompass the full query-response space, leading to suboptimal convergence and performance limitations, which our IPO framework aims to address by iteratively refining the alignment process and enhancing the coverage and quality of the policy optimization.
[23] demonstrates that reward models may exhibit superior generalization capabilities compared to policies in terms of sample complexity, meaning that, given a fixed number of samples, reward models achieve higher preference classification accuracy. To leverage this insight, we first train a critic model capable of performing both pairwise ranking and pointwise scoring. During the training process, no additional human feedback is queried; instead, the trained critic model is utilized to label the samples generated by the model in subsequent training iterations. However, since the model is initially poorly aligned with human feedback, the generated results often suffer from low quality, leading to inefficient sampling. This results in collected samples that may not adequately approximate the target distribution, ultimately causing performance degradation when training on these samples. To address this issue and enhance algorithmic efficiency, we propose an iterative preference optimization (IPO) framework, which progressively refines the model’s alignment with human preferences, enabling a more effective and gradual achievement of the final optimization goal.
In RLHF, the optimization goal is to maximize the expected reward given the condition $c$, while increasing the KL-divergence as a reward penalty to keep the optimized model close to the reference distribution.
@@
where the hyperparameter $\beta$ controls regularization and is usually a fixed value.
In [57], a loss functional is defined in terms of $\pi_{\theta}(\cdot|c)$, expressed as
$\mathbb{E}_{x_{0}\sim\pi(x_{0}|c)}\Big{[}-r(x_{0},c)-\beta\log\frac{\pi_{ref}(x_{0}|c)}{\pi_{\theta}(x_{0}|c)}\Big{]}=\\ \beta\mathbb{KL}\Big{(}\pi_{\theta}(x_{0}|c)\Big{\|}\pi_{ref}(x_{0}|c)\exp\Big{(}\frac{1}{\beta}r(x_{0},c)\Big{)}\Big{)},$ (2)
the value that minimizes the loss functional is $\pi^{*}(x_{0}|c)\propto\pi_{0}(x_{0}|c)\exp\Big{(}\frac{1}{\beta}r(x_{0},c)\Big{)}$. This is consistent with the results of DPO, which directly transforms the optimization objective into training $\pi^{*}(x_{0}|c)$.
The conclusion drawn is that the reinforcement learning process, when combined with the maximum expected reward under KL regularization, aims to approximate $\pi^{*}(x_{0}|c)$ during the learning process. However, in practical scenarios, obtaining unlimited offline datasets to achieve this goal is infeasible, and the model often achieves only limited improvement and constrained generalization when trained on finite and suboptimal offline datasets. While it is necessary to utilize the Critic model to annotate the results generated by the model, the efficiency and effectiveness of data generation often lead to suboptimal outcomes in a single training iteration, highlighting the need for iterative refinement to enhance performance and generalization capabilities.
Inspired by [31], we introduce an iterative preference optimization method that, rather than directly approximating $\pi_{\theta}$ by $\pi_{\text{ref}}\exp\left(\frac{1}{\beta}r(x_{0},c)\right)$—a challenging task—divides the process into multiple stages by examining a series of intermediate distributions. By approximating $\pi_{\text{ref}}\exp\left(\frac{1}{\beta_{i-1}}r\right)$ with $\pi_{\text{ref}}\exp\left(\frac{1}{\beta_{i-1}}r\right)$ at each iteration, we achieve stepwise improvements, making the optimization process more tractable and effective.
$\pi_{ref}^{(0)}=\pi_{ref},$
$\pi_{ref}^{(1)}=\pi_{ref}^{(0)}\exp\Big{(}\frac{1}{\beta_{1}}r\Big{)},$
$\pi_{ref}^{(2)}=\pi_{ref}^{(1)}\exp\Big{(}\frac{1}{\beta_{2}}r\Big{)},$
$\vdots$
$\pi^{*}=\pi_{ref}^{(N)}\exp\Big{(}\frac{1}{\beta_{2}}r\Big{)}.$
To validate the effectiveness and generalizability of the proposed approach, we introduce two distinct offline preference alignment methodologies: Diffusion-DPO, which leverages pairwise ranking data to align models with human preferences, and Diffusion-KTO, which utilizes pointwise scoring data to optimize alignment objectives, both designed to enhance the performance and adaptability of the framework across diverse scenarios.
Diffusin-DPO Objective
Diffusion-DPO [45] extends DPO [36] by introducing inverse decomposition to replace $\pi_{\theta}$ and $\pi_{ref}$. The forward process $q(x_{1:T}|x_{0})$ serves as an approximation for the reverse process $p_{\theta}(x_{1:T}|x_{0})$. Ultimately, the objective function is formulated after these adjustments. By redefining data likelihood in the context of diffusion models, Diffusion-DPO introduces a straightforward and effective loss formulation, enabling robust and efficient preference optimization.
$L_{diffusion-dpo}(\theta)=\\ -\mathbb{E}_{(x^{w}_{0},x^{l}_{0})\sim D,t\sim\mathcal{U}(0,T),x^{w}_{t}\sim q(x^{w}_{t}|x^{w}_{0}),x^{l}_{t}\sim q(x^{l}_{t}|x^{l}_{0})}\\ \log\sigma\left(-\beta T\omega(\lambda_{t})\left(\right.\right.\\ \|\epsilon^{w}-\epsilon_{\theta}(x_{t}^{w},t)\|^{2}_{2}-\|\epsilon^{w}-\epsilon_{\text{ref}}(x_{t}^{w},t)\|^{2}_{2}\\ \left.\left.-\left(\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\|^{2}_{2}-\|\epsilon^{l}-\epsilon_{\text{ref}}(x_{t}^{l},t)\|^{2}_{2}\right)\right)\right)$ (3)
where $x^{w}$ is the chosen sample, $x^{l}$ is the rejected sample, $\omega(\lambda_{t})$ a weighting function [15, 18, 36]).
Diffusin-KTO Objective
Inspired by prospect theory, KTO proposes the concept of human-aware loss functions by maximizing the utility function and defining reference points. KTO does not require preference data and can directly use data marked with binary signals to train the algorithm, making it more sensitive to negative samples. Instead of optimizing the expected reward, Diffusion-KTO adopts a nonlinear utility function that calculates the utility of an action based on its value $Q(a,s)$ relative to a reference point $Q_{\text{ref}}$.
@@
$L_{diffusion-kto}(\theta)=\underset{\pi_{\theta}}{\text{max}}\mathbb{E}_{x_{0}\sim\mathcal{D},t\sim\text{Uniform([0,T])}}\\ [U(w(x_{0})(\beta\log\frac{\pi_{\theta}(x_{t-1}|x_{t})}{\pi_{\text{ref}}(x_{t-1}|x_{t})}-Q_{\text{ref}}))]$ (4)
where $w(x_{0})=\pm 1$, $x_{0}$ represents a state that is considered either desirable or undesirable. The quantity $Q_{ref}$ is introduced as $\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}(a|s)||\pi_{\text{ref}}(a|s)]$, which measures the Kullback-Leibler divergence between the policy $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$.
Iterative Preference Optimization (IPO) framework, each iteration training consists of three key components: video generation, critic, and preference learning. The implementation details will be introduced in the next section.
4 Implementation details
In this section, we present our iterative preference optimization framework, along with its implementation details and pipeline. An overview of the algorithm is illustrated in Fig. 1. Specifically, the data collection and annotation process are described in detail in Section 4.1, while Section 4.2 elaborates on the training methodology for the Critic model. Finally, Section 4.3 introduces a comprehensive framework for aligning human feedback, encompassing the training of DPO on paired data and KTO on single data.
4.1 Datasets Annotation Pipeline
Video-text collection
Our annotated dataset is constructed using the open-source model CogVideoX-2B [52]. The process begins by creating a diverse set of prompts for video generation, encompassing various character types, actions, scenes (both realistic and abstract), animals, plants, vehicles, and video styles. These elements are randomly combined and refined into coherent prompts using the Qwen2.5 [51] large language model. To ensure diversity and avoid the generation of repetitive sentences, we fine-tune the model’s output by adjusting key parameters such as temperature, top-k, and top-p. Using these prompts, the video generation model produces three distinct variants for each video by applying different random seeds. Throughout the training process, we iteratively regenerate data using an optimized strategy to progressively expand and enhance our preference-aligned dataset.
Human Annotation
When learning to align human feedback, we focus on three dimensions, referring to the text-to-image evaluation work Evalalign[43]: consistency between text and video, video fidelity, and additional motion smoothness. To capture human preferences, we employ two annotation methods: Pairwise ranking and Pointwise scoring.
For pairwise ranking, annotators are presented with two videos generated from the same prompt and instructed to evaluate them based on predefined criteria. During the evaluation, annotators are required to: (1) provide detailed natural language feedback, (2) identify specific quality issues in each video (or confirm their absence), and (3) articulate the reasoning behind their ranking decisions. These annotated comparisons are then systematically organized to construct the training samples:
For pointwise scoring, each video is evaluated across three dimensions, with scores categorized into three levels: good, average, and poor. Similar to the ranking process, annotators are required to provide detailed justifications for their scores, ensuring transparency and consistency in the evaluation process. These annotated comparisons are then systematically organized to construct the training samples:
Each dataset is annotated by at least three individuals. The final annotation result is determined through voting, selecting the annotation with the highest level of consistency among the reviewers.
4.2 Critic model training
To develop a comprehensive model capable of handling both paired data sorting and individual video scoring tasks, we fine-tuned the pre-trained Vila model [25], leveraging its exceptional capabilities in image and video instruction following, comprehension, and multi-image processing, which make it particularly suitable for training our Critic model. During training, we implemented key modifications, including restructuring the prompt format to incorporate detailed reasoning chains before point-by-point scoring or pairwise ranking, and constructing explicit chains of thought to enhance the model’s decision-making process. This approach enables the model to provide comprehensive rationales prior to delivering its final output, improving both the accuracy of results and the quality of natural language feedback.
We initialized the model using the Vila 13B/40B pre-trained checkpoint and fine-tuned it for 5 epochs on our labeled dataset with a learning rate of 2e-5, distributed across 32 GPUs with a batch size of 4 per GPU. To mitigate inconsistencies from scaling and center cropping, we implemented pad training. For better video content understanding, we extracted 16 uniformly sampled frames from each video sequence, keeping other hyperparameters aligned with Vila’s defaults. This setup optimizes both computational efficiency and model performance, ensuring strong video-text alignment.
4.3 Iterative training pipeline
Iterative preference optimization provides substantial advantages over offline learning methods. The iterative learning framework we propose is a versatile and general-purpose approach, capable of seamlessly integrating various offline learning algorithms into an iterative setting, thereby achieving significant performance enhancements. Specifically, we investigate two distinct offline preference alignment methods: the KTO algorithm, which utilizes pointwise scoring, and the DPO algorithm, which is based on pairwise ranking.
Iterative DPO
In the iterative preference optimization process employing Direct Preference Optimization (DPO), the system follows a structured workflow: during each iteration, the current optimal policy model generates multiple video outputs for a given prompt. These generated videos are subsequently evaluated and ranked by a critic model based on predefined quality metrics. The ranked outputs then serve as training data for preference alignment, where the policy model is fine-tuned to better align with the desired preferences. This cyclic process of generation, evaluation, and optimization continues iteratively, progressively enhancing the model’s performance through successive refinement cycles.
  • Video Generation. We use CogvideoX-2B as the initial checkpoint, and generate at least three random videos for the same prompt by setting different seeds to to form a dataset $\mathcal{V}$ = $\{(x_{k},y_{k_{1}},y_{k_{2}},y_{k_{3}},...)\}_{k=1}^{N}$, where $x_{k}$ represents the prompt and $y_{k}$ represents the corresponding video generated by the model.
  • Pairwise Ranking. To enhance ranking reliability and mitigate potential biases in the Critic model, we implement a robust pairwise comparison protocol with position swapping during inference: for each video pair, we conduct two reciprocal evaluations by alternating their presentation order, retaining only consistent results as valid entries for the preference dataset. Formally, we construct the pairwise dataset as $\mathcal{D}=\{(x_{k},y_{k_{w}},y_{k_{l}})\}_{k=1}^{N}$, where $y_{k_{w}}$ and $y_{k_{l}}$ denote the preferred (winning) and dispreferred (losing) samples, respectively. This rigorous validation process ensures higher data quality and reduces positional bias in the collected preference pairs.
  • Preference learning. During training, we employ a selective sampling strategy to construct high-quality preference pairs by extracting only the highest-ranked and lowest-ranked samples from the critic model’s evaluations, forming the final paired data $(y^{w},y^{l})$ while deliberately excluding intermediate-ranked samples to ensure clearer preference distinctions. For optimization, we utilize the diffusion-DPO loss as the primary objective function and introduce two auxiliary negative log-likelihood (NLL) loss terms following  [33]: the first term, weighted by 0.2, regularizes the preferred samples, while the second term, scaled by 0.1, anchors the model to real video data distributions. This dual regularization strategy preserves the structural integrity and format of generated outputs while preventing undesirable degradation in the log-probability of selected responses, as demonstrated in  [32],  [11], and  [33].
    $\mathcal{L}(\theta)$ $=L_{diffusion-dpo}(\theta)$ (5)

### v3

@@ Abstract
Video foundation models have achieved significant advancement with the help of network upgrade as well as model scale-up. However, they are still hard to meet requirements of applications due to unsatisfied generation quality. To solve this problem, we propose to align video foundation models with human preferences from the perspective of post-training in this paper. Consequently, we introduce an Iterative Preference Optimization strategy to enhance generated video quality by incorporating human feedback. Specifically, IPO exploits a critic model to justify video generations for pairwise ranking as in Direct Preference Optimization or point-wise scoring as in Kahneman-Tversky Optimization. Given this, IPO optimizes video foundation models with guidance of signals from preference feedback, which helps improve generated video quality in subject consistency, motion smoothness and aesthetic quality, etc. In addition, IPO incorporates the critic model with the multi-modality large language model, which enables it to automatically assign preference labels without need of retraining or relabeling. In this way, IPO can efficiently perform multi-round preference optimization in an iterative manner, without the need of tediously manual labeling. Comprehensive experiments demonstrate that the proposed IPO can effectively improve the video generation quality of a pretrained model and help a model with only 2B parameters surpass the one with 5B parameters. Besides, IPO achieves new state-of-the-art performance on VBench benchmark.The source code and model checkpoints, along with implementation details and resources, are openly available at this GitHub repository.
Figure 1: Comparison of (a) traditional single-round preference optimization and (b) the proposed Iterative Preference Optimization for aligning video foundation model with human feedback. IPO introduces a critic model to automatically labeling video data without need of tedious manual efforts, making iterative update possible in an efficient way.
Figure 2: The overview of our proposed Iterative Preference Optimization framework. IPO consists of three major parts: (a) Human Preference Dataset. It is used to train a critic model. This dataset only requires one-round manual annotation without need of relabeling in the iterative framework; (b) Critic Model. It is learned to automatically annotate generated videos with preference labels, which help eliminate tedious manual efforts for multi-round optimization; (c) Iterative Optimization. It incorporates the critical model to iteratively optimize video foundation model. In this way, IPO can efficiently tune the base model to align with human preference and enhance generation ability in aspects of subject consistency, motion smoothness and aesthetic quality, etc. Best viewed in 2x zoom.
@@ 1 Introduction
Thus, we propose to align Text-to-Video models (T2Vs) with human preference from the perspective of post-training in this paper. In particular, we present the Iterative Preference Optimization (IPO) framework for enhancing T2Vs. IPO introduces preference priors for finetuning T2Vs in the form of reinforcement learning, which considers subject consistency, motion smoothness and aesthetic quality, etc. This leads to improved video generations that are well aligned with human preference. In addition, as show in Fig. 1, difference from prior single-round methods, IPO adopts a multi-round optimization strategy to reinforce T2Vs in an iterative manner, which strengths the models continuously. In this way, IPO can effectively improve T2Vs without need of large-scale dataset and expensive supervised finetuning.
@@ 2 Related Work
Text-to-video generation Recent advances in video generation have been largely driven by diffusion models [11, 31, 32, 46], achieving remarkable improvements in diversity, fidelity, and quality [4, 6, 13, 20, 29, 38, 40]. Scaling model size and training data further enhance performance [12, 30, 39]. In Text-to-Video (T2V) generation, pre-training from scratch requires substantial computational resources and large-scale datasets [8, 22, 16, 2, 15]. Current methods [1, 37, 9] extend text-to-image models by incorporating spatial and temporal attention modules, often using joint image-video training [23], with [9] proposing a plug-and-play temporal module for personalized image models. State-of-the-art pixel quality and temporal consistency have been achieved by [47, 3], while diffusion models based on the DiT structure [2, 43, 3] enhance spatiotemporal coherence and scalability, highlighting the potential of Transformer-based architectures. Despite of their success, they still cannot meet requirements in real scenarios due to lack of alignment with human preference.
Post-training for human preference alignment Post-training is a common technique exploited in Large Language Models (LLMs)s [24, 28, 50]. Reinforcement Learning from Human Feedback (RLHF) [24] learns a reward function from comparison data to align the policy model. Rank Responses for Human Feedback (RRHF) [45] extends supervised fine-tuning, while Sequence Likelihood Calibration (SLiC) [49] leverages offline RL data. Direct Preference Optimization (DPO) [27] improves stability and efficiency by directly optimizing human preference data, and Kahneman-Tversky Optimization (KTO) [7] aligns models via a prospect-theoretic approach. Rejection Sampling Optimization (RSO) [21] enhances alignment using statistical rejection sampling. Prior works have introduced this technique into enhancement of image foundation models, such as [5, 25, 41]. Diffusion-DPO [36] extends DPO for diffusion models, optimizing model likelihood, while Diffusion-KTO [18] maximizes expected human utility without costly pairwise preference data. The usage of post-training in video generation is still rare. Current works mainly focus on the design of reward models. [44] introduced a temporal decaying reward (TAR) mechanism, prioritizing central frames for stable and coherent video generation. Meanwhile, [17] optimized ordinary differential equation (ODE) solvers by exploring energy function designs, enhancing T2V training through motion prior extraction. VADER [26] fine-tunes video diffusion models using reward model gradients and pre-trained vision models, achieving strong generalization even for unseen cues, thereby improving T2V adaptability and performance. However, the above strategies can only achieve minor improvements and not fully explore the potential of post-training.
Figure 3: The distribution statistics of Human Preference Dataset on prompts in (a) and (b) as well as scoring/ranking in (c) and (d).
@@ 3 METHOD
The proposed Iterative Preference Optimization framework consists of three major parts: (1) A human preference dataset for training critic model; (2) A critic model for automatically labeling generated videos; (3) An iterative optimization process for enhancing the base text-to-video models. Fig. 2 illustrates the whole pipeline of IPO. We will explain details of each part in the following sections.
3.1 Human Preference Dataset
Video-text Collection
The generated video dataset is created using the CogVideoX-2B model [43]. We generate diverse prompts covering various character types, which are then combined and refined using the Qwen2.5 [42]. To increase diversity, we adjust parameters like temperature, top-k, and top-p. As shown in Fig. 3, the dataset includes over ten categories, such as Human, Architecture, Vehicles, and Animals, as well as more than ten different actions of human and animal. For each prompt, the video generation model produces four variants with different random seeds. The dataset is iteratively expanded through an optimized regeneration strategy to improve preference alignment.
Human Annotation When aligning human feedback, we focus on three dimensions, inspired by the text-to-image evaluation work Evalalign [34]: consistency between text and video, video faithfulness, and additional motion smoothness. To capture human preferences, we employ two annotation methods: Pairwise Ranking and Pointwise Scoring. For pairwise ranking, annotators are shown two videos generated from the same prompt and asked to evaluate them based on predefined criteria. They are required to: (1) provide detailed natural language feedback, and (2) identify specific quality issues in each video. The annotated comparisons are then organized to create training samples.
For pointwise scoring, each video is assessed across three dimensions, with scores categorized into three levels: Good, Normal, and Poor. Similar to the ranking process, annotators provide detailed justifications for their scores to ensure transparency and consistency. The annotated evaluations are then organized to create training samples.
Each dataset is annotated by at least three individuals. The final annotation is determined through voting, selecting the one with the highest consistency among reviewers.
3.2 Critic Model Training
Large Visual-Language Models (VLMs) exhibit strong generalization across diverse tasks due to pre-training on extensive multimodal data. Their capability to capture complex text-image relationships makes them well-suited for critic modeling, where alignment with human preferences is crucial. Leveraging VLMs as critic models enhances video generation by improving preference predictions through their robust generalization ability.
Fig. 2 illustrates our supervised fine-tuning process. The training starts by passing video frames through an image encoder to extract image tokens. These tokens, together with the textual caption, question, and answer, are then fed into a pre-trained Visual-Language Model (VLM). Formally, each training sample consists of a triplet: multimodal input $M$, which includes both the video and its corresponding textual description; question $Q$; and answer $A$.
The optimization objective is derived from the autoregressive loss function commonly used in LLM training, but applied specifically to the answer as:
$L(\phi)=-\sum^{N}_{i=1}\log p(A_{i}\mid Q,M,A_{<i};\phi),$ (1)
where $N$ is the length of the ground-truth answer. This loss encourages the model to get answers that align with provided questions and multimodal context. Depending on the method, the model receives different inputs. In the pointwise case, it evaluates a single video-text pair. In the pairwise case, it compares two pairs to determine which aligns better with the text. The optimization focuses on individual quality in pointwise and relative comparison in pairwise.
3.3 T2V Alignment
3.3.1 Alignment Optimization
In Reinforcement Learning with Human Feedback (RLHF), the primary objective is to optimize the model to maximize the expected reward given context $c$. This is achieved by incorporating a regularization term that penalizes deviations from a reference distribution, ensuring the optimized model remains consistent with prior knowledge or previously trained models. The optimization objective is as:
@@ 3.3.1 Alignment Optimization
where $r(x_{0},c)$ denotes the reward associated with a state $x_{0}$ given context $c$, $\pi_{\theta}(x_{0}|c)$ is the policy of the model parameterized by $\theta$, and $\pi_{ref}(x_{0}|c)$ is the reference policy. The hyperparameter $\beta$ controls the strength of the regularization, typically set as a fixed constant. This objective function aims to balance reward maximization with regularization, which prevents the model from straying too far from the reference distribution, ensuring stability and generalization [48]. And the loss function for the policy $\pi_{\theta}(\cdot|c)$ is formulated in terms of the expected value over the reward and the KL-divergence penalty as following:
$\mathbb{E}_{x_{0}\sim\pi_{\theta}(x_{0}|c)}\Big{[}-r(x_{0},c)-\beta\log\frac{\pi_{ref}(x_{0}|c)}{\pi_{\theta}(x_{0}|c)}\Big{]}=\\ \beta\mathbb{KL}\left(\pi_{\theta}(x_{0}|c)\Big{\|}\pi_{ref}(x_{0}|c)\exp\left(\frac{1}{\beta}r(x_{0},c)\right)\right),$ (3)
Minimizing this loss leads to optimal policy $\pi^{*}(x_{0}|c)$ as:
$\pi^{*}(x_{0}|c)\propto\pi_{0}(x_{0}|c)\exp\left(\frac{1}{\beta}r(x_{0},c)\right),$ (4)
This aligns with Direct Preference Optimization (DPO) [27] that translates the optimization goal into training the optimal policy $\pi^{*}(x_{0}|c)$. The key advantage of this formulation is that it incorporates both reward maximization and regularization into a unified framework, thereby enabling robust learning. To improve the model’s alignment with human feedback and output quality, we incorporate real-world data into the alignment training process. By using authentic, human-generated examples, the model learns to reduce hallucinations and better align with human preferences. This also enhances the model’s robustness and generalization, preventing training divergence and ensuring reliable performance in diverse real-world tasks.
3.3.2 Pointwise Optimization
For pointwise preference alignment, we adopt Diffusion-KTO [7]. Unlike traditional preference-based methods, KTO optimizes a nonlinear utility function using binary-labeled data, making it particularly sensitive to negative samples. This approach avoids the need for explicit preference pairs and introduces human-aware loss functions.
In Diffusion-KTO, the utility of an action is computed relative to a reference point $Q_{\text{ref}}$, defined as the KL divergence between the learned policy $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$. The optimization objective is:
@@ 3.3.2 Pointwise Optimization
$L_{\text{kto}}(\theta)=\underset{\pi_{\theta}}{\text{max}}\mathbb{E}_{x_{0}\sim\mathcal{D},t\sim\text{Uniform([0,T])}}\\ [U(w(x_{0})(\beta\log\frac{\pi_{\theta}(x_{t-1}|x_{t})}{\pi_{\text{ref}}(x_{t-1}|x_{t})}-Q_{\text{ref}}))],$ (5)
where $w(x_{0})=\pm 1$ indicates whether a state $x_{0}$ is desirable or undesirable. The reference point $Q_{\text{ref}}$ is defined as:
$Q_{\text{ref}}=\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}(a|s)||\pi_{\text{ref}}(a|s)].$ (6)
This adaptive reference mechanism allows Diffusion-KTO to calibrate the policy toward human-aligned decision-making. To optimize the policy, we use the loss function from Eq. 5. A challenge in KTO training is estimating the average reward $Q$, which becomes unstable with small batch sizes. To mitigate this, we set the batch size as 128 for stable $Q$ estimation. To further improve stability and generalization, we incorporate real-world data from the VidGen-1M dataset [33]. We add a negative log-likelihood (NLL) loss term on real samples to regularize the model and prevent divergence from natural video distributions. The final training objective is:
$\mathcal{L}(\theta)=L_{\text{kto}}(\theta)+\lambda\cdot\mathbb{E}_{(\textbf{{y}},x)\sim\mathcal{D}^{\mathrm{real}}}\bigl{[}-\log p_{\theta}(y|x)\bigr{]},$ (7)
where $\lambda$ controls the weight of real data regularization. This hybrid approach balances the model’s utility maximization with adherence to real-world video distributions, enhancing its alignment with human preferences.
3.3.3 Pairwise Optimization
For pairwise preference alignment, we adopt Diffusion-DPO [36], an extension of Direct Preference Optimization (DPO) [27] to diffusion models. Diffusion-DPO utilizes inverse decomposition to replace the standard policy $\pi_{\theta}$ and reference policy $\pi_{\text{ref}}$, enabling efficient preference optimization within the diffusion framework.
In Diffusion-DPO, the forward process $q(x_{1:T}|x_{0})$ approximates the reverse process $p_{\theta}(x_{1:T}|x_{0})$, redefining data likelihood in diffusion models. This leads to a loss function that facilitates robust preference optimization. The objective function for Diffusion-DPO is as:
$L_{\text{dpo}}(\theta)=-\mathbb{E}_{(x^{w}_{0},x^{l}_{0})\sim D,t\sim\mathcal{U}(0,T),x^{w}_{t}\sim q(x^{w}_{t}|x^{w}_{0}),x^{l}_{t}\sim q(x^{l}_{t}|x^{l}_{0})}\\ \log\sigma\left(-\beta T\omega(\lambda_{t})\left(\|\epsilon^{w}-\epsilon_{\theta}(x_{t}^{w},t)\|^{2}_{2}-\|\epsilon^{w}-\epsilon_{\text{ref}}(x_{t}^{w},t)\|^{2}_{2}\right.\right.\\ -\left.\left(\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\|^{2}_{2}-\|\epsilon^{l}-\epsilon_{\text{ref}}(x_{t}^{l},t)\|^{2}_{2}\right)\right),$ (8)
where $x^{w}$ denotes the preferred sample, $x^{l}$ the rejected sample, and $\omega(\lambda_{t})$ is a weighting function [11, 14, 27].
During training, we use a selective sampling strategy to construct high-quality preference pairs by selecting the highest-ranked (preferred) and lowest-ranked (rejected) samples based on critic model evaluations. This ensures clearer preference distinctions and enhances learning. These pairs $(y^{w},y^{l})$ form the final training data.
The optimization objective combines the Diffusion-DPO loss with two auxiliary negative log-likelihood (NLL) terms. The first, weighted by $\lambda_{1}{=}0.2$, regularizes the preferred samples to align with human preferences. The second, weighted by $\lambda_{2}{=}0.1$, anchors the model to real video distributions to prevent overfitting. The final loss is:
$\mathcal{L}(\theta)$ $=L_{\text{dpo}}(\theta)$ (9)